How many molecules are present in 176 grams of Vitamin C (chemical formula C6H8O6)?
Answer #1

Answer

Molar mass of Vitamin C = (12 x 6) + (1 x 8) + (16 x 6) = 72 + 8 + 96 = 176 g/mol

So, no. of moles of vitamin c in 176 g = 176g/(176 g/mol) = 1 mol